Emma launches a ball off of a platform of height h with velocity v0 at an angle θ above the horizontal. A wall with a ball-sized hole a hight of H above the floor, with H>h, is a horizontal distance L= 16.0 m away from Emma. The difference in height between H and h is one fifth of L, the square of the magnitude of the velocity, v02, is five-fourths of the product gL.

Determine the two values of θ that ensure the ball passes through the hole.

Hint: It is possible to use the quadratic formula to solve any quadratic equation that is equal to zero, even when the variable itself is a function.
Answer #1

hole ginen L=16m V² - 9,8x 16 Vo = 12.5 m/s H: h= 16 & L=16m condition is that pays Condition is that ball passes through hol4²-42 = 2gs 0-12 sin20 = -2 X 9.84 16 = = R sin20= 2X 9.8 x 16 I displacement ish-b) 9.8 x 16 s sin?o= 2 = sino = + 1 2 3 whe